OpenStudy (anonymous):

solve be completing the square: x^2+7x-4=0

OpenStudy (anonymous):

Add and substract LHS by 49/4 \[x^{2}-7x+\frac{ 49 }{ 4 }-\frac{ 49 }{ 4 }-4=0\] \[[x-\frac{ 7 }{ 2 }]^{2}-\frac{ 65 }{ 4 }=0\] \[[x-\frac{ 7 }{ 4 }]^2=\frac{ 65 }{ 4 }\] further take square root and solve further....